# Generative Consistency Models for Estimation of Kinetic Parametric Image Posteriors in Total-Body PET

### v1

@@ Abstract
Dynamic total body positron emission tomography (TB-PET) makes it feasible to measure the kinetics of all organs in the body simultaneously which may lead to important applications in multi-organ disease and systems physiology. Since whole-body kinetics are highly heterogeneous with variable signal-to-noise ratios, parametric images should ideally comprise not only point estimates but also measures of posterior statistical uncertainty. However, standard Bayesian techniques, such as Markov chain Monte Carlo (MCMC), are computationally prohibitive at the total body scale. We introduce a generative consistency model (CM) that generates samples from the posterior distributions of the kinetic model parameters given measured time-activity curves and arterial input function. CM is able to collapse the hundreds of iterations required by standard diffusion models into just 3 denoising steps. When trained on 500,000 physiologically realistic two-tissue compartment model simulations, the CM produces similar accuracy to MCMC (median absolute percent error $5\%$; median K-L divergence <0.5<0.5) but is more than five orders of magnitude faster. CM produces more reliable $\mathit{K_{i}}$ images than the Patlak method by avoiding the assumption of irreversibility, while also offering valuable information on statistical uncertainty of parameter estimates and the underlying model. The proposed framework removes the computational barrier to routine, fully Bayesian parametric imaging in TB-PET and is readily extensible to other tracers and compartment models.
@@ 1 Introduction
Dynamic positron emission tomography (PET) provides time-resolved measurements of a radiotracer distribution in the body, to which compartmental models can be fit for the estimation of physiological parameters, such as the net influx rate constant, $K_{i}$. These quantitative parameters enable the testing of hypotheses in research studies, while in clinical studies they may provide additional discriminatory power beyond that of the semi-quantitative standardised uptake value (SUV) [1]. Common parameter estimation techniques include nonlinear least squares and linear graphical analyses, such as the Patlak plot, which are more computationally efficient [2, 3, 4, 5]. The recent introduction of total-body and long axial field of view ($\geq 1$ metre) PET scanners makes it possible to acquire dynamic data from all major organs of the body simultaneously. However, this comes at the computational cost of estimating model parameters over more than four million voxels [6]. It also introduces new challenges, in particular the heterogeneity of kinetics throughout the human body, calling into question the appropriateness of assuming a single kinetic model for all organs and tissues. As well as physiological variability, signal-to-noise ratios are highly variable throughout the body making it increasingly important to compute, not only point estimates of model parameters, but also measures of their statistical uncertainty.
@@
Bayesian inference addresses this need by computing the posterior distributions, $p(\mathbf{\theta}|y)$, of the kinetic model parameters, $\mathbf{\theta}$, given a time-activity curve (TAC), $y$. Markov chain Monte Carlo (MCMC) is the reference method for generating asymptotically unbiased posteriors [7, 8]. However, MCMC requires an analytical expression for the likelihood function based on an assumed error distribution, which is not reliably known at the voxel level when using non-linear reconstruction methods such as EM, as well as a long burn-in phase and careful assessment of convergence. Even when convergence is reached, MCMC samplers are notoriously slow and are not easy to parallelize. In dynamic total-body PET (TB-PET) studies, where millions of voxels require independent MCMC chains (a single study may contain 4.4 million voxels), these characteristics render MCMC impractical [9]. Likelihood-free alternatives, such as approximate Bayesian computation (ABC), improve the situation by bypassing the need to compute the likelihood and are more amenable to parallelisation, but they are still computationally highly inefficient [10]. Variational Bayesian methods offer a faster, deterministic alternative, but their approximations can lead to biased parameter estimates and inaccurate uncertainty quantification [11]. As a result, the application of Bayesian computation methods for posterior probability estimation in PET is largely confined to the analysis of regions of interest.
@@
Deep generative models have recently emerged as computationally efficient alternatives to MCMC. Conditional variational auto-encoders (CVAEs), generative adversarial networks (GANs) and denoising diffusion probabilistic models (DDPMs) have all been applied to medical imaging to estimate Bayesian posteriors, reducing inference time while maintaining reasonable performance compared with MCMC [12, 13, 14]. Nonetheless, these approaches have limitations. CVAEs may severely underestimate posterior variance [11], GANs suffer from mode collapse [15], and diffusion models require hundreds to thousands of reverse steps [16], which scales poorly to TB-PET image volumes. Moreover, existing studies have focused mainly on brain imaging data [14] and the applicability of generative models to dynamic TB-PET, with its broader range of kinetics and greater inter-organ heterogeneity, has not been demonstrated.
@@
The present work introduces a conditional consistency model (CM) framework for posterior estimation in dynamic TB-PET. The CM reformulates diffusion sampling as a short schedule of denoising operations and has been shown to retain the theoretical guarantees of score-based diffusion (SBD) while collapsing the sampling path to three steps [17]. We adapt this idea to PET compartmental modelling by training a CM to learn a direct mapping from pure Gaussian noise samples to kinetic parameter samples, conditioned on both the TAC and the arterial input function (AIF). Trained on a large ensemble of physiologically realistic simulations, the proposed method produces voxel-wise posterior samples for the given kinetic model within a practical time frame, without sacrificing accuracy. We evaluate the performance of the CM by comparison with the reference method MCMC, as well as ABC and common generative deep learning approaches. We then illustrate a practical application of the CM method to a dynamic total body [18F]FDG PET study, and we show how results from CM can be used to perform voxel-level model selection tasks to account for kinetic heterogeneity and to compute parametric images of posterior uncertainty in estimating $K_{i}$.
@@ 2.1 Consistency Models
[…]
@@ 2.1.1 Model Architectures
The building block of the proposed conditional consistency model is a 1D U‐Net (Fig. 2) that processes two inputs: the noisy kinetic parameter vector $\mathbf{\theta}_{t}\in\mathbb{R}^{\mathbf{\theta}_{1\times n}}$ and the concatenated dynamic measurements $y\in\mathbb{R}^{y_{1\times m}}$ consisting of a TAC and the corresponding AIF, together with a scalar noise level $t$. First, $\mathbf{\theta}_{t}$ is linearly transformed to a vector of size 32 for the 1D convolutions, while $y$ is embedded through a dense layer. Simultaneously, $t$ is embedded via Gaussian Fourier features followed by a linear projection. The two resulting embeddings are combined via element-wise SiLU activation [19]. The SiLU is smooth and non-monotonic, with a non-zero derivative even for negative inputs, which often leads to faster convergence and improved gradient flow compared to ReLU or other piecewise‐linear activations. This combined embedding vector is then injected through a dense layer at every level of the U‐Net to condition all convolutional blocks on both $y$ and $t$. The encoder (downsampling) path consists of four successive 1D convolutional blocks with kernel size 3. The first block maps $1\to 32$ channels with stride 1, the second maps $32\to 64$ channels with stride 2, the third maps $64\to 128$ channels with stride 2, and the fourth maps $128\to 256$ channels with stride 2. After each convolution, the network adds the dense projection of the combined embedding, applies group normalization, and then applies the SiLU activation. Residual (skip) connections carry the output of each encoder block forward into the matching decoder stage. The decoder (upsampling) path mirrors this structure using transposed 1D convolutions: $256\to 128$, $128\to 64$, $64\to 32$, and finally $32\to 1$ channels. At each upsampling stage, the decoder concatenates the corresponding encoder feature map, injects the same combined embedding via a dense layer, applies group normalization, and uses SiLU activation to restore spatial resolution. A final dense layer projects the length of 32 feature map back to the original dimension $\mathbf{\theta}_{1\times n}$, yielding a denoised $\hat{\mathbf{\theta}}_{0}$.
@@ 2.2.1 Two-Tissue Compartment Model
$C_{t}(t)=(1-V_{b})(C_{f}(t)+C_{m}(t))+V_{b}C_{b}(t).$
To account for measurement noise, a Gaussian noise model is introduced [21], and the measured PET time series is assumed to be,
@@ 2.2.2 Training Datasets
The kinetic parameter vectors $(K_{1},k_{2},k_{3},k_{4},V_{b})$ were sampled uniformly from ranges reported in the PET literature for FDG in a wide variety of organs and tissues [22, 23, 24]. Specifically, we chose
@@
$k_{4}$ $0].\displaystyle\in[-10,010]\ \mathrm{min^{-1}},V_{b}\in[03,020$].
which together span the physiological variability reported in human PET studies. We allowed $k_{4}$ to be negative to allow the possibility of negative values in the posterior which facilitates statistical analysis to test whether $k_{4}$ is non‐zero, as will become apparent. Note that a negative $k_{4}$ is physiologically plausible as it can be interpreted as effectively augmenting the forward transfer rate $k_{3}$. The AIF $C_{b}(t)$ was simulated using the Feng model [25],
@@
where the seven parameters $(\beta_{1},\beta_{2},\beta_{3},\kappa_{1},\kappa_{2},\kappa_{3},t_{0})$ were drawn uniformly from ranges covering the full variability observed in eight subjects,
@@
This parameterization ensured that the training set included diverse input function shapes and delays, enabling the model to learn from and generalize across realistic physiological conditions. Noise was added to the noise-free simulation $C_{t}(t)$ via the noise model described in Section 2.2.1 with noise level $l$ randomly drawn from $3$ to $7$. The range of noise levels was derived from clinical dynamic FDG-PET data from eight subjects. The noisy measured value is then
@@ 2.2.3 Training Details
The 1D U-Net was trained for 60,000 epochs with a batch size of 200,000 on the synthetic training dataset. We optimized the student model with AdamW with fixed learning rate $0.0016$. A separate teacher model was trained via an EMA of the student parameters with a fixed decay rate of 0.99. At each iteration, two adjacent noise levels drawn uniformly from a fixed schedule of $N_{k}=50$ steps, were used to compute the consistency loss between the student and teacher outputs, and the teacher parameters were updated. For comparison, we tested an adaptive schedule in which $N_{k}$ and EMA $\mu$ started at small values and were gradually increased to the preset maximum $N_{k}$ and $\mu$ over the course of training, allowing coarser noise discretization early on and finer resolution later. We also tested various choices for $N_{k}$ (from 2 to 200). Finally, we found that the fixed 50-step noise schedule and fixed EMA $\mu$ yielded the best validation performance. We trained all models on a cluster of Nvidia A100 GPUs.
@@ 2.3.1 Reference Method: MCMC
Posterior ground truth was generated with the No-U-Turn Sampler (NUTS) implemented in PyMC [26, 27]. NUTS is an adaptive Hamiltonian Monte Carlo algorithm that eliminates the need to hand-tune trajectory length by automatically terminating each leap-frog path when it begins to double back on itself, thereby maintaining detailed balance while exploring high-dimensional parameter spaces efficiently. For every simulated or voxel-wise analysis we ran four independent chains, each with 10,000 warm-up iterations used for dual averaging of the step size and mass-matrix adaptation, followed by 200,000 posterior draws per chain. A target acceptance probability of $0.9$ was retained, and the sampler’s automatic jittering of initial values ensured robust exploration even for poorly scaled parameters. Convergence was confirmed when all monitored parameters satisfied the Gelman–Rubin convergence statistic $\hat{R}<1.01$ [28] and the effective sample size exceeded 40,000. Runs that failed these criteria were repeated with a longer warm-up or more sampling steps.
@@ 2.3.2 Compared Methods
ABC provides a likelihood‑free reference rooted in forward simulation: kinetic parameter vectors are drawn from the prior, TACs are simulated with the compartmental model, and a draw is retained only when the simulated TAC lies within a preset distance from the measured one [10]. The accepted draws form an empirical posterior but, because proposals are not guided toward high‑probability regions, the acceptance rate decays rapidly with dimensionality. In our study, we instantiate ABC for dynamic PET using PET-ABC [29].
The DDPM benchmarks posterior estimation by learning the reverse of a 1000‑step Gaussian noise process that maps a standard normal prior to the conditional posterior $p(\theta|y)$. During training, the network reconstructs the added noise at each diffusion level, optimizing a mean squared error that has a closed‑form connection to variational inference [16]. We implemented the denoiser with the same 1‑D U‑Net backbone, depth and parameter count as CM to ensure capacity parity. Inference follows the original reverse chain to generate samples, incurring 1000 evaluations per draw.
SBD uses a continuous stochastic differential equation (SDE) whose forward form gradually perturbs data toward Gaussian noise, while a learned score function drives the reverse SDE back to the posterior [18]. Training minimizes a denoising score matching loss that directly estimates the score $\nabla_{\theta_{t}}\log p_{t}(\theta_{t}\mid y)$ at different noise levels. A U‑Net identical to the one used in our CM was used. Sampling integrates the reverse SDE with an Euler–Maruyama solver for 1000 time steps.
For CVAE, two fully connected encoders map $(\theta,y)$ and $y$ alone to the mean and log‑variance of a latent Gaussian. Reparameterised noise is then concatenated with $y$ and passed through a fully connected decoder to reconstruct a kinetic parameter sample. For more details about CVAE, readers are directed to [12]. We reused the CVAE-dual-decoder described in this reference. The three networks (i.e., two encoders and one decoder) were up‑sized so the number of learnable parameters matches the number of parameters in the U-Net used in CM.
The GAN trains a generator $G_{\mathbf{\theta}}$ to transform a Gaussian noise vector concatenated with the 70‑element TAC–AIF vector into plausible parameter samples, and a critic $D_{\psi}$ that enforces the Wasserstein distance with a gradient‑penalty regulariser. After convergence, posterior draws require only a single forward pass through $G_{\mathbf{\theta}}$. Readers are directed to [30] for more details. The generator and critic adopt encoder–decoder U‑Nets whose width gives the total number of parameters comparable to CM and the diffusion baselines.

### v2

@@ Abstract
Dynamic total body positron emission tomography (TB-PET) makes it feasible to measure the kinetics of the tracer in all organs of the body simultaneously which may lead to important applications in multi-organ disease and systems physiology. Since whole-body kinetics are highly heterogeneous with variable signal-to-noise ratios, parametric images should ideally comprise not only point estimates but also measures of posterior statistical uncertainty. However, standard Bayesian techniques, such as Markov chain Monte Carlo (MCMC), are computationally prohibitive at the total body scale. We introduce a generative consistency model (CM) that generates samples from the posterior distributions of the kinetic model parameters given measured time-activity curves and arterial input function. CM is able to collapse the hundreds of iterations required by standard diffusion models into just 3 denoising steps. The CM was evaluated using physiologically realistic simulations and an application to a subject’s dynamic [18F]FDG TB-PET dataset analyzed with a standard single-input two-tissue compartment model. When trained on 500,000 physiologically realistic two-tissue compartment model simulations, the CM produces similar accuracy to MCMC (median absolute percent error $5\%$; median K-L divergence <0.5<0.5) but is more than five orders of magnitude faster. CM produces more reliable $\mathit{K_{i}}$ images than the Patlak method by avoiding the assumption of irreversibility, while also offering valuable information on statistical uncertainty of parameter estimates and the underlying model. The proposed framework removes the computational barrier to routine, fully Bayesian parametric imaging in TB-PET and is readily extensible to other tracers and compartment models.
@@ 1 Introduction
Dynamic positron emission tomography (PET) provides time-resolved measurements of a radiotracer distribution in the body, to which compartmental models can be fit for the estimation of physiological parameters, such as the net influx rate constant, $K_{i}$. These quantitative parameters enable the testing of hypotheses in research studies, while in clinical studies they may provide additional discriminatory power beyond that of the semi-quantitative standardised uptake value (SUV) [1]. Common parameter estimation techniques include nonlinear least squares and linear graphical analyses, such as the Patlak plot, which are more computationally efficient [2, 3, 4, 5]. The recent introduction of total-body and long axial field of view ($\geq 1$ metre) PET scanners makes it possible to acquire dynamic data from all major organs of the body simultaneously. Current efforts in total-body parametric imaging have primarily utilized graphical methods like Patlak to manage the computational cost of estimating parameters over millions of voxels [6, 7]. However, these graphical approaches often suffer from bias due to simplifying assumptions (e.g., irreversibility), while standard nonlinear least squares (NLS) remains computationally prohibitive for such large volumes. Emerging deep learning approaches have begun to address these challenges [8, 9, 10], yet they predominantly focus on non-Bayesian estimation, leaving the critical aspect of uncertainty quantification largely unexplored. Therefore, the gap is significant because total-body imaging introduces substantial kinetic heterogeneity, calling into question the appropriateness of assuming a single kinetic model for all organs and tissues. Furthermore, signal-to-noise ratios are highly variable throughout the body, making it increasingly important to compute not only point estimates of kinetic parameters but also measures of their statistical uncertainty to assess reliability.
@@
Bayesian inference addresses this need by computing the posterior distributions, $p(\mathbf{\theta}|y)$, of the kinetic model parameters, $\mathbf{\theta}$, given a time-activity curve (TAC), $y$. Markov chain Monte Carlo (MCMC) is the reference method for generating asymptotically unbiased posteriors [11, 12]. However, MCMC requires an analytical expression for the likelihood function based on an assumed noise distribution, which is not reliably known at the voxel level when using non-linear reconstruction methods such as expectation maximization, as well as a long burn-in phase and careful assessment of convergence. Even when convergence is reached, MCMC samplers are slow and are not easy to parallelize. In dynamic total-body PET (TB-PET) studies, where millions of voxels require independent MCMC chains (a single study may contain millions of voxels), these characteristics make MCMC impractical [13]. Likelihood-free alternatives, such as approximate Bayesian computation (ABC), improve the situation by bypassing the need to compute the likelihood, but they are still computationally highly inefficient [14]. Variational Bayesian methods offer a faster alternative, but their approximations can lead to biased parameter estimates [15]. As a result, the application of Bayesian computation methods for posterior probability estimation in PET is largely confined to the analysis of regions of interest.
@@
Deep generative models have recently emerged as computationally efficient alternatives to MCMC. Conditional variational auto-encoders (CVAEs), generative adversarial networks (GANs) and denoising diffusion probabilistic models (DDPMs) have all been applied to medical imaging to estimate posteriors [16, 17, 18]. Nonetheless, these approaches have limitations. CVAEs underestimate posterior variance [15], GANs suffer from mode collapse [19], and diffusion models require hundreds to thousands of reverse steps [20], which scales poorly to TB-PET image volumes. Moreover, existing studies have focused mainly on ROI-based analysis of the brain [18] and the application of generative models to dynamic TB-PET, with its broader range of kinetics and greater inter-organ heterogeneity, has not been demonstrated.
@@
The present work introduces a conditional consistency model (CM) for posterior estimation in dynamic TB-PET. The CM reformulates diffusion sampling as a short schedule of denoising operations and has been shown to retain the theoretical guarantees of score-based diffusion (SBD) while collapsing the sampling path to one step [21]. We adapt this idea to PET compartmental modelling by training a CM to learn a direct mapping from pure Gaussian noise samples to kinetic parameter samples, conditioned on both the TAC and the arterial input function (AIF). Trained on a large ensemble of physiologically realistic simulations, the proposed method produces voxel-wise posterior samples for the given kinetic model within a practical time frame, without sacrificing accuracy. We evaluate the performance of the CM by comparison with the reference method MCMC, as well as ABC and common generative deep learning approaches. We then illustrate a practical application of the CM method to a dynamic total body [18F]FDG PET study, and we show how results from CM can be used to perform voxel-level model selection tasks to account for kinetic heterogeneity and to compute parametric images of posterior uncertainty in estimating $K_{i}$.
@@ 2.1 Consistency Models
The core of this training algorithm is to teach the model $f_{\phi}$ to produce the same output in one forward pass at a higher noise level as an exponential moving average (EMA) “teacher” model $f_{\phi^{-}}$ would produce at a slightly lower noise level. Each iteration begins by sampling a pair $(\mathbf{\theta},y)$ from the dataset $\mathcal{D}$ and drawing a noise index $n$ uniformly from $\{1,\dots,N\}$. A Gaussian vector $z$ is scaled by $t_{n+1}$ and $t_{n}$ to create two noisy versions: $\mathbf{\theta}+t_{n+1}z$ (more corrupted) and $\mathbf{\theta}+t_{n}z$ (slightly less corrupted). The student network $f_{\phi}$ is evaluated on the more corrupted input, while the teacher network is evaluated on the slightly cleaner input. The distance
@@ 2.1.1 Model Architectures
The building block of the proposed conditional consistency model is a 1D U‐Net (Fig. 2) that processes two inputs: the noisy kinetic parameter vector $\mathbf{\theta}_{t}\in\mathbb{R}^{\mathbf{\theta}_{1\times n}}$ and the concatenated dynamic measurements $y\in\mathbb{R}^{y_{1\times m}}$ consisting of a TAC and the corresponding AIF, together with a scalar noise level $t$. First, $\mathbf{\theta}_{t}$ is transformed to a vector of size 32 (a dimension determined empirically to balance representational capacity with computational efficiency) for the following 1D convolutions, while $y$ is embedded through a dense layer. Simultaneously, $t$ is embedded via Gaussian Fourier features followed by a linear projection. The two resulting embeddings are combined via element-wise SiLU activation [23]. The SiLU is smooth and non-monotonic, with a non-zero derivative even for negative inputs, which often leads to faster convergence and improved gradient flow compared to ReLU or other piecewise‐linear activations. This combined embedding vector is then injected through a dense layer at every level of the U‐Net to condition all convolutional blocks on both $y$ and $t$. The encoder (downsampling) path consists of four successive 1D convolutional blocks with kernel size 3. The first block maps $1\to 32$ channels with stride 1, the second maps $32\to 64$ channels with stride 2, the third maps $64\to 128$ channels with stride 2, and the fourth maps $128\to 256$ channels with stride 2. After each convolution, the network adds the dense projection of the combined embedding, applies group normalization, and then applies the SiLU activation. Residual (skip) connections carry the output of each encoder block forward into the matching decoder stage. The decoder (upsampling) path mirrors this structure using transposed 1D convolutions: $256\to 128$, $128\to 64$, $64\to 32$, and finally $32\to 1$ channels. At each upsampling stage, the decoder concatenates the corresponding encoder feature map, injects the same combined embedding via a dense layer, applies group normalization, and uses SiLU activation to restore spatial resolution. A final dense layer projects the length of 32 feature map back to the original dimension $\mathbf{\theta}_{1\times n}$, yielding a denoised $\hat{\mathbf{\theta}}_{0}$.
@@ 2.2.1 Two-Tissue Compartment Model
$C_{t}(t)=(1-V_{b})(C_{f}(t)+C_{m}(t))+V_{b}C_{wb}(t).$
where $C_{wb}(t)$ represents the whole-blood radioactivity concentration. In this study, both $C_{b}(t)$ and $C_{wb}(t)$ were approximated using the image-derived input function extracted from the ascending aorta. To account for measurement noise, a Gaussian noise model is introduced [25], and the measured PET time series is assumed to be,
@@ 2.2.2 Training Datasets
The kinetic parameter vectors $(K_{1},k_{2},k_{3},k_{4},V_{b})$ were sampled uniformly from ranges reported in the PET literature for FDG in a wide variety of organs and tissues [26, 27]. Specifically, we chose
@@
$k_{4}$ $0],\displaystyle\in[-10,010]\ \mathrm{min^{-1}},V_{b}\in[03,020$],
where these intervals were found to encapsulate the vast majority of physiological tissues. We allowed $k_{4}$ to take negative values during network training. This allows for a robust estimation of the parameter distribution when the true value lies near the boundary of zero. In other words, exposing the model to negative $k_{4}$ values ensures that network predictions around $k_{4}=0$ are within the training range and are reliable [28], enabling the robust model selection analysis described in Section 3.4.
The AIF $C_{b}(t)$ was simulated using the Feng model [29],
@@
[…]
Note that by explicitly varying $t_{0}$ during training, the model learns features invariant to tracer arrival delays, effectively performing an implicit delay correction during inference. This parameterization ensured that the training set included diverse input function shapes (representing variations in injection speed/dispersion) and delays, enabling the model to learn from and generalize across realistic physiological conditions. Noise was added to the noise-free simulation $C_{t}(t)$ via the noise model described in Section 2.2.1 with noise level $l$ randomly drawn from $3$ to $7$. The range of noise levels was also derived from the clinical dynamic FDG PET data from the eight subjects.
The noisy measured value is then
@@ 2.2.3 Training Details
The 1D U-Net was trained using the AdamW optimizer with a batch size of 200,000. Training was scheduled for a maximum of 80,000 epochs. The validation loss was observed to stabilize and plateau after 150,000 iterations. Consequently, the final model weights were selected from epoch 60,200 (approx. 150,500 iterations), as this checkpoint corresponded to the minimum validation error. A separate teacher model was trained via an EMA of the student parameters with a fixed decay rate of 0.99. At each iteration, two adjacent noise levels drawn uniformly from a fixed schedule of $N_{k}=50$ steps, were used to compute the consistency loss between the student and teacher outputs, and the teacher parameters were updated. For comparison, we tested an adaptive schedule in which $N_{k}$ and EMA $\mu$ started at small values and were gradually increased to the preset maximum $N_{k}$ and $\mu$ over the course of training, allowing coarser noise discretization early on and finer resolution later. We also tested various choices for $N_{k}$ (from 2 to 200). Finally, we found that the fixed 50-step noise schedule and fixed EMA $\mu$ yielded the best validation performance. We trained and evaluated all deep learning models on a single node equipped with 4 NVIDIA A100 GPUs (80GB VRAM) using data parallelisation.
@@ 2.3.1 Reference Method: MCMC
Posterior ground truth was generated with the adaptive Hamiltonian Monte Carlo algorithm implemented in PyMC [30, 31]. For inference we ran four independent chains, each with 10,000 warm-up iterations, followed by 200,000 posterior draws per chain. A target acceptance probability of $0.9$ was retained, and the sampler’s automatic jittering of initial values ensured robust exploration even for poorly scaled parameters. To ensure the MCMC samples served as a reliable ground truth for the posterior distribution, strict convergence criteria were enforced. Convergence was confirmed only when all kinetic parameters satisfied the Gelman-Rubin convergence statistic $\hat{R}<1.01$ [32] and the effective sample size exceeded 40,000. Chains failing these strict diagnostics were discarded and re-sampled with increased warm-up phases to guarantee that the MCMC reference represented the true stationary distribution.
@@ 2.3.2 Compared Methods
ABC provides a likelihood‑free reference rooted in forward simulation: kinetic parameter vectors are drawn from the prior, TACs are simulated with the compartmental model, and a draw is retained only when the simulated TAC lies within a preset distance from the measured one [14]. The accepted draws form an empirical posterior but, because proposals are not guided toward high‑probability regions, the acceptance rate decays rapidly with dimensionality. In our study, we instantiate ABC for dynamic PET using PET-ABC [33]. The DDPM benchmarks posterior estimation by learning the reverse of a 1000‑step Gaussian noise process that maps a standard normal prior to the conditional posterior $p(\theta|y)$. During training, the network reconstructs the added noise at each diffusion level, optimizing a mean squared error that has a closed‑form connection to variational inference [20]. We implemented the denoiser with the same 1‑D U‑Net backbone, depth and parameter count as CM to ensure capacity parity. Inference follows the original reverse chain to generate samples, incurring 1000 evaluations per draw. SBD uses a continuous stochastic differential equation (SDE) whose forward form gradually perturbs data toward Gaussian noise, while a learned score function drives the reverse SDE back to the posterior [22]. Training minimizes a denoising score matching loss that directly estimates the score $\nabla_{\theta_{t}}\log p_{t}(\theta_{t}\mid y)$ at different noise levels. A U‑Net identical to the one used in our CM was used. Sampling integrates the reverse SDE with an Euler–Maruyama solver for 1000 time steps. For CVAE, two fully connected encoders map $(\theta,y)$ and $y$ alone to the mean and log‑variance of a latent Gaussian. Reparameterised noise is then concatenated with $y$ and passed through a fully connected decoder to reconstruct a kinetic parameter sample. For more details about CVAE, readers are directed to [16]. We reused the CVAE-dual-decoder described in this reference. The three networks (i.e., two encoders and one decoder) were up‑sized so the number of learnable parameters matches the number of parameters in the U-Net used in CM. The GAN trains a generator $G_{\mathbf{\theta}}$ to transform a Gaussian noise vector concatenated with the 70‑element TAC–AIF vector into plausible parameter samples, and a critic $D_{\psi}$ that enforces the Wasserstein distance with a gradient‑penalty regulariser. After convergence, posterior draws require only a single forward pass through $G_{\mathbf{\theta}}$. Readers are directed to [34] for more details. The generator and critic adopt encoder–decoder U‑Nets whose width gives the total number of parameters comparable to CM and the diffusion baselines.